OpenStudy (anonymous):

Slope-intercept form to coordinates: Intersection of y = 2x+13 and y + -1/2x + 13 Intersection of 2x - y = 17 and y = -1/2x + 13

OpenStudy (anonymous):

anyone?

OpenStudy (aum):

"y + -1/2x + 13" Is that supposed to be: y = -1/2x + 13 (= in the place of the first +) ?

OpenStudy (anonymous):

Yeah.